# LLM-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

### v1

@@ Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting. However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series. To address this, we propose LLM-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment. Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts. Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information. Finally, the transformed time series from the LLMs are projected to obtain the forecasts. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that LLM-Prompt is a powerful framework for time series forecasting.
1 Introduction
@@
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions [1]. Over the past five years, neural networks with diverse architectures—including linear models [2, 3], convolutional neural networks [4, 5], recurrent neural networks [6], and Transformers [7]—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings [8, 9]. On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance [10, 11].
@@
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting [12, 13, 14]. However, due to the modality gap between LLMs and time series data, knowledge transfer is required. As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers [16]; b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering [9, 15]; c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e. a hybrid architecture of the above two strategies [17, 18].
@@
Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist. On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance [19, 20]. For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series [9]. Overall, a unified prompt paradigm remains absent. On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities [21, 22]. Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
To address the aforementioned technical challenges, this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed LLM-Prompt. The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture. Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data. Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation. Finally, the LLMs’ outputs are projected to obtain time series forecasting results. Experimental results demonstrate that LLM-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners. The main contributions of this study are summarized as follows:
  • We propose the LLM-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
@@
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-fusion of temporal data and textual information, enhancing feature representation.
@@
  • Extensive experiments on 7 benchmark datasets and 3 carbon emission datasets demonstrate that LLM-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
2 Related Work
@@
Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures. Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures. iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms [23]. PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization [24]. Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series [25]. Other variants include Autoformer [26], which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer [27], which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting. DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency [2]. TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities [3]. CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years. SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions [28]. WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs [29]. RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently. Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting [6]. This method dynamically selects critical temporal features to capture long-term dependencies. However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios [32]. As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs. Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations. Prompt-based LLMs convert numerical time series into textual information. Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations. However, challenges remain in activating LLMs’ forecasting capabilities. Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts [17, 20]; LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data [8]; Time-LLM is limited to constructing hard prompts using only statistical features of temporal data [19]. Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap [30, 31].
3 Methodology
3.1 Framework Overview
As illustrated in Figure 2, LLM-Prompt comprises three key modules: dual-path prompting, cross-modal data fusion (including semantic space embedding and cross-modal alignment), and time series forecasting.
@@ 3.1 Framework Overview
Figure 2: LLM-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
3.2 Dual-Path Prompting
@@
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series [17, 20]. First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
@@
Prompt Pool Construction. Given prompt keys $K\in{R^{P\times D}}$ and prompt pool $V\in{R^{P\times L\times D}}$, where P is the pool size, L is the soft prompt length, and D is the embedding dimension. The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs. Notably, the prompt pool is updated during network iterations. Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
@@
Figure 3: Hard prompt example
Hard Prompts. Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs. However, a well-considered unified paradigm remains absent. Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches [19, 21]. Experimental practice indicates this limits performance gains. To address this, we propose a unified hard prompt paradigm. As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features. Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities. Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
3.3 Cross-Modal Data Fusion
Semantic Space Embedding. This component primarily involves embedding time series data and hard prompts [21, 31, 22]. For soft prompts, they are generated directly in the embedding space and thus require no embedding.
$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
3.4 Time Series Forecasting
As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM. Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$. GPT-2 is used by default unless otherwise specified, and subsequent experiments also investigate the impact of different LLM architectures on forecasting performance [33].
4 Experiments
To validate the effectiveness of LLM-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting. Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM [19], TimeCMA [21]; (2) Transformer-based: Autoformer [26], Crossformer [35], iTransformer[23], PatchTST [36]; (3) CNN-based: TimeNet [5]; (4) RNN-based: Mamba[6]; (5) MLP-based: DLinear [2].
4.1 Long-term Forecasting
Setups. We conduct experiments on 9 datasets, including 6 classical benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, exchange-rate, weather) and 3 carbon emission datasets (munichCarbon, chinaCarbon, worldCarbon). The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that LLM-Prompt outperforms all baselines in most cases. LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations. Compared to other LLMs-based approaches, we observe that LLM-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA. This is attributed to our integration of both soft/hard prompts and cross-modal alignment in LLM-Prompt.
Table 1: Long-term forecasting results. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@ 4.1 Long-term Forecasting
4.2 Few-shot Forecasting
Results. The results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. LLM-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
Table 2: Few-shot learning on 10% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@ 4.2 Few-shot Forecasting
Table 3: Few-shot learning on 5% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@
4.3 Zero-shot Forecasting
Table 4: Zero-shot learning on carbon emission datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Results. Brief results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other deep learning approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting. In the $ETTh1\rightarrow ETTm1$ setting, LLM-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
Table 5: Zero-shot learning on ETT datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@ 4.3 Zero-shot Forecasting
We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global). Results demonstrate that LLM-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
4.4 Ablation Study
Ablations on Hard Prompt. We further investigate how different types of information within hard prompts affect LLM-Prompt’s performance. As shown in Figure 3, we compare LLM-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration. As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃performance degradation.
Table 6: Ablations on different components. Red: the best.
Methods LLM-Prompt w/o Soft Prompt w/o Hard Prompt w/o Cross-Modal Alignment
@@ 4.4 Ablation Study
Table 7: Ablations on different hard prompt configurations. Red: the best.
Methods LLM-Prompt w/o Historical Data w/o Statistical Data
@@
Ablations on Soft Prompt. Additionally, we investigate the impact of soft prompt pool size and soft prompt length on LLM-Prompt’s performance. As shown in Figure 5, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon. This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
@@
5 Conclusion
This paper proposes LLM-Prompt to unlock the potential of LLMs in time series forecasting. The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks. To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features. Extensive experiments on multiple real-world datasets validate the effectiveness of LLM-Prompt in time series forecasting.
Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks. Ours future work will explore integrating Chain-of-Thought techniques to unlock LLMs’ predictive potential or developing foundation models tailored for forecasting tasks.

### v2

# TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

@@ Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting. However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series. To address this, we propose TSF-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment. Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts. Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information. Finally, the transformed time series from the LLMs are projected to obtain the forecasts. Furthermore, we focus on applying time series forecasting to carbon emissions, aiming to provide a modest contribution to global carbon neutrality. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that TSF-Prompt is a powerful framework for time series forecasting.
Codehttps://github.com/SanMuGuo/Time-Prompt
Introduction
@@
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions (Wang et al. 2024b). Over the past five years, neural networks with diverse architectures—including linear models (Zeng et al. 2023; Wang et al. 2024a), convolutional neural networks (Wang et al. 2023; Wu et al. 2023), recurrent neural networks (Gu and Dao 2023), and Transformers (Vaswani et al. 2017)—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings (Gruver et al. 2023; Liu et al. 2024b). On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al. 2024; Ni et al. 2025).
@@
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting (Shyalika et al. 2024; Jin et al. 2023b; Zhang et al. 2024). However, due to the modality gap between LLMs and time series data, knowledge transfer is required. As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers (Zhou et al. 2023); b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al. 2024b; Xue and Salim 2023); c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e. a hybrid architecture of the above two strategies (Cao et al. 2023; Jia et al. 2024).
@@
Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist. On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance (Jin et al. 2023a; Pan et al. 2024). For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series (Liu et al. 2024b). Overall, a unified prompt paradigm remains absent. On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities (Liu et al. 2024a, 2025). Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
To address the aforementioned technical challenges, inspired by recent research (Liu et al. 2024a; Cao et al. 2023; Jin et al. 2023a), this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed TSF-Prompt. The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture. Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data. Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation. Finally, the LLMs’ outputs are projected to obtain time series forecasting results. Experimental results demonstrate that TSF-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners.
Furthermore, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals. The main contributions of this study are summarized as follows:
  • We propose the TSF-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
@@
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, enhancing feature representation.
@@
  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that TSF-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
Related Work
@@
Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures. Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures. iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms (Liu et al. 2023). PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization (Nie et al. 2023). Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series (Liu et al. 2022b). Other variants include Autoformer (Wu et al. 2021), which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer (Zhou et al. 2022), which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
@@
Figure 2: TSF-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting. DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency (Zeng et al. 2023). TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities (Wang et al. 2024a). CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years. SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions (Liu et al. 2022a). WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs (Van Den Oord et al. 2016). RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently. Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting (Gu and Dao 2023). This method dynamically selects critical temporal features to capture long-term dependencies. However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al. 2024c). As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs. Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations. Prompt-based LLMs convert numerical time series into textual information. Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations. However, challenges remain in activating LLMs’ forecasting capabilities. Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts (Cao et al. 2023; Pan et al. 2024); LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data (Gruver et al. 2023); Time-LLM is limited to constructing hard prompts using only statistical features of temporal data (Jin et al. 2023a). Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap (Kowsher et al. 2024; Hu et al. 2025).
Methodology
Framework Overview
As illustrated in Figure 2, TSF-Prompt comprises four components: dual-path prompting, semantic space embedding, cross-modal alignment, and time series forecasting. First, hard and soft prompts are constructed based on temporal data. Subsequently, semantic space embedding is performed through the pre-trained LLM and patch reprogramming module. Since soft prompts are inherently generated within this space, no additional embedding is required. Next, cross-modal alignment is executed. Finally, the fused information is fed into the pre-trained LLM and projected through output layers to generate forecasting results.
Dual-Path Prompting
@@
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series (Cao et al. 2023; Pan et al. 2024). First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
@@
Prompt Pool Construction. Given prompt keys $K\in{R^{P\times D}}$ and prompt pool $V\in{R^{P\times L\times D}}$, where P is the pool size, L is the soft prompt length, and D is the embedding dimension. The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs. Notably, the prompt pool is updated during backpropagation. Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
@@
Figure 3: Hard prompt example on chinaCarbon dataset.
Hard Prompts. Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs. However, a well-considered unified paradigm remains absent. Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches (Jin et al. 2023a; Liu et al. 2024a). Experimental practice indicates this limits performance gains. To address this, we propose a unified hard prompt paradigm. As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features. Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities. Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
Table 1: Long-term forecasting results. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@
Table 2: Few-shot learning on 10% training data. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@
Cross-Modal Data Fusion
Semantic Space Embedding. This component primarily involves embedding time series data and hard prompts (Liu et al. 2024a; Hu et al. 2025; Liu et al. 2025). For soft prompts, they are generated directly in the embedding space and thus require no embedding.
$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Time Series Forecasting
As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM. Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$. GPT-2 is used by default unless otherwise specified (Radford et al. 2019).
Experiments
To validate the effectiveness of TSF-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting. Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE). All experiments are conducted on V100 GPUs. Additional details (e.g., datasets, SOTA methods, full results) are provided in the appendix.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al. 2023a), TimeCMA (Liu et al. 2024a); (2) Transformer-based: Autoformer (Wu et al. 2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al. 2023), PatchTST (Nie et al. 2022); (3) CNN-based: TimeNet (Wu et al. 2023); (4) RNN-based: Mamba(Gu and Dao 2023); (5) MLP-based: DLinear (Zeng et al. 2023).
Long-term Forecasting
Setups. We conduct experiments on 9 datasets, including 6 classical benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, exchange-rate, weather) and 3 carbon emission datasets (munichCarbon, chinaCarbon, worldCarbon). The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others.
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that TSF-Prompt outperforms all baselines in most cases. LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations. Compared to other LLMs-based approaches, we observe that TSF-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA. This is attributed to our integration of both soft/hard prompts and cross-modal alignment in TSF-Prompt.
Table 3: Few-shot learning on 5% training data. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@
Table 4: Zero-shot learning on ETT datasets. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
@@
Few-shot Forecasting
Results. The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. TSF-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
Zero-shot Forecasting
Results. Results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting. In the $ETTh1\rightarrow ETTm1$ setting, TSF-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global). Results demonstrate that TSF-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
Table 5: Zero-shot learning on carbon emission datasets. Bold: the best, Underline: the second best.
Methods TSF-Prompt Time-LLM iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Table 6: Ablations on different components. Bold: the best.
Methods TSF-Prompt w/o Soft Prompt w/o Hard Prompt w/o Alignment
@@
Table 7: Ablations on different hard prompt configurations.
Methods TSF-Prompt w/o Historical Data w/o Statistical Data
@@
Ablation Study
Ablations on Hard Prompt. We further investigate how different types of information within hard prompts affect TSF-Prompt’s performance. As shown in Figure 3, we compare TSF-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration. As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃% performance degradation.
Ablations on Soft Prompt. Additionally, we investigate the impact of soft prompt pool size and soft prompt length on TSF-Prompt’s performance. As shown in Figure 4, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon. This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
Conclusion
This paper proposes TSF-Prompt to unlock the potential of LLMs in time series forecasting. The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks. To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features. Extensive experiments on multiple real-world datasets validate the effectiveness of TSF-Prompt in time series forecasting.
Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks. Our future work will explore developing foundation models specifically tailored for forecasting tasks, or further demonstrate the predictive capabilities of LLMs through techniques such as chain-of-thought reasoning, fine-tuning, and reinforcement learning.